# Tube-NeRF: Efficient Imitation Learning of Visuomotor Policies from MPC using Tube-Guided Data Augmentation and NeRFs

### 23 Nov 2023 v1

@@ Abstract
Imitation learning (IL) can train computationally-efficient sensorimotor policies from a resource-intensive Model Predictive Controller (MPC), but it often requires many samples, leading to long training times or limited robustness. To address these issues, we combine IL with a variant of robust MPC that accounts for process and sensing uncertainties, and we design a data augmentation (DA) strategy that enables efficient learning of vision-based policies. The proposed DA method, named Tube-NeRF, leverages Neural Radiance Fields (NeRFs) to generate novel synthetic images, and uses properties of the robust MPC (the tube) to select relevant views and to efficiently compute the corresponding actions. We tailor our approach to the task of localization and trajectory tracking on a multirotor, by learning a visuomotor policy that generates control actions using images from the onboard camera as only source of horizontal position. Our evaluations numerically demonstrate learning of a robust visuomotor policy with an $80$-fold increase in demonstration efficiency and a $50\%$ reduction in training time over current IL methods. Additionally, our policies successfully transfer to a real multirotor, achieving accurate localization and low tracking errors despite large disturbances, with an onboard inference time of only $1.5$ ms. Video: https://youtu.be/_W5z33ZK1m4
@@ I INTRODUCTION
Imitation learning (IL) [1, 2, 3] has been extensively employed to train sensorimotor neural network (NN) policies for computationally-efficient onboard sensing, planning and control on mobile robots. Training data is provided by computationally expensive model-based planners/controllers [4, 5, 6], acting as expert demonstrators. The resulting policies produce commands from raw sensory inputs, bypassing the computational cost of control (e.g., solving large optimization problems in model predictive control (MPC) [7]) and sensing (e.g., localization).
@@
However, one of the fundamental limitations of existing IL methods employed to train sensorimotor policies (e.g., Behavior Cloning (BC)) [2], DAgger [3]) is the overall number of demonstrations that must be collected from the model-based algorithms. This inefficiency is rooted in an issue known as covariate shift [3], i.e. when the training data distribution differs significantly from the one at deployment, and is worsened by uncertainties (e.g., in the robot/environment dynamics [8, 9]).
@@
Figure 1: Efficient visuomotor policy learning approach: Tube-NeRF collects a real-world demonstration using an output-feedback variant of robust tube MPC, which accounts for process and sensing uncertainties through its tube cross-section $\mathbb{Z}$. Then, it generates a Neural Radiance Field (NeRF) of the environment from the collected images $\mathcal{I}_{t}$, and uses the NeRF for data augmentation. The tube’s cross-section guides the selection of additional views, $\mathcal{I}_{t}^{+}$, while the corresponding actions are efficiently computed via the ancillary controller, an integral component of the tube MPC framework. An image rendered from the NeRF of the MIT Highbay flight space is shown in the background.
Strategies based on matching the uncertainties between the training and deployment domains (e.g., Domain Randomization (DR) [10, 11, 12, 4, 5]) improve the robustness of the learned policy, but create challenges in data- and computational-efficiency, as they require applying a large set of perturbations throughout multiple demonstrations, repeatedly querying a computationally-expensive expert.
@@
Leveraging high-fidelity simulators on powerful computers avoids challenging data-collection on the real robot, simplifying DR, but it introduces sim-to-real gaps that are especially challenging when learning visuomotor policies, i.e., that directly use raw pixels as input. For this reason, sensorimotor policies trained in simulation often leverage as input easy-to-transfer visual abstractions, such as feature tracks [5], depth maps [13, 14], intermediate layers of a convolutional NN (CNN) [15], or a learned latent space [16]. However, all these abstractions discard information that may instead benefit task performance.
@@
Data augmentation (DA) approaches that augment demonstrations with extra images and stabilizing actions improve robustness and sample efficiency of IL [17, 18, 19, 20, 21]. However, they (i) rely on handcrafted heuristics for the selection of extra images and the generation of corresponding actions, (ii) do not explicitly account for the effects of uncertainties when generating the extra data [17, 18, 19, 20, 21], and (iii) often leverage ad-hoc image acquisition setups for DA [18, 19]. As a consequence, their real-world deployment has been mainly focused on tasks in 2D (e.g., steering a Dubins car [17, 19]).
@@
In this work, summarized in Figure 1, we propose Tube- Neural Radiance Field (NeRF), a new DA framework that enables efficient learning of robust visuomotor policies from MPC and that overcomes the aforementioned limitations in DA. Tube-NeRF extends our previous DA strategy for IL [9], where we showed that a robust variant of MPC, robust tube model predictive controller (RTMPC) [22], can be used to efficiently and systematically generate extra data for sample-efficient learning of motor control policies (e.g., actions from full state). The new work presented herein enables efficient learning of policies that directly use vision and other sensors as input, relaxing the constraining assumption in our previous work [9] that full-state information is available at deployment.
@@
Our new strategy first collects robust task-demonstrations that account for the effects of process and sensing uncertainties via an output-feedback variant of RTMPC [23]. Then, we employ a photorealistic representation of the environment, based on a NeRF, to generate novel views for DA, using the tube of the controller to guide the selection of the extra novel views and sensorial inputs, and using the ancillary controller to efficiently compute the corresponding actions. We further enhance DA by organizing collected real-world observations in a database and employing the tube to generate queries. This enables usage of real-world data for DA, reducing the sim-to-real gap when only synthetic images are used. Lastly, we adapt our approach to a multirotor, training a visuomotor policy for robust trajectory tracking and localization using onboard camera images and additional measurements of altitude, orientation, and velocity. The generated policy relies solely on images to obtain information on the robot’s horizontal position, which is a challenging task due to (1) its high speed (up to $3.5$ m/s), (2) varying altitude, (3) aggressive roll/pitch changes, (4) the sparsity of visual features in our flight space, and (5) the presence of a safety net that moves due to the down-wash of the propellers and that produces semi-transparent visual features above the ground.
This Letter is an evolution of our previous conference paper [24], where we demonstrated the capabilities of our framework in simulation. In this extension, (i) we present the first-ever real-time deployment of our approach, demonstrating (in more than $30$ flights) successful agile trajectory tracking with policies that use onboard camera images to infer the horizontal position of a multirotor despite aggressive 3D motion, and subject to a variety of sensing and dynamics disturbances. Our policy has an average inference time of only $1.5$ ms onboard a small GPU (Nvidia Jetson TX2) and is deployed at $200$ Hz. To achieve these results, we (ii) leverage a NeRF instead of 3D meshes (as in [24]) for photo-realistic novel view-synthesis for DA; (iii) modify our DA strategy to include real-world images during DA, further reducing the sim-to-real gap in the deployment of our policies, (iv) employ a computationally-efficient but capable visuomotor NN architecture for onboard deployment, and (v) use randomization in image space to increase robustness to visual changes in the environment.
Contributions:
  1. 1)
@@
    A new DA strategy enabling efficient (demonstrations, training time) learning of a sensorimotor policy from MPC. The policy generates actions using raw images and other measurements, instead of the full-state estimate [9], and is robust in the real world to a variety of uncertainties. Our approach is grounded in the output feedback RTMPC framework theory, unlike previous DA methods that rely on handcrafted heuristics, and uses a NeRF to generate images.
  2. 2)
@@
    A procedure to apply our methodology for tracking and localization on a multirotor, using onboard images from a fisheye camera and altitude, attitude and velocity data.
  3. 3)
    Numerical and experimental evaluations of a policy learned from a single demonstration that runs onboard in $1.5$ ms on average under a variety of different uncertainties (wind, a slung-load causing visual disturbances, sensing noise).
@@ II RELATED WORKS
Method Domain of training data Policy directly uses images Avoids specialized data-collection equipment Demonstration-efficient Explicitly accounts for uncertainties Avoids Hand-Crafted Heuristics for Data Augm. Real-world deployment (2D/3D, Domain)
@@
Novel view synthesis with NeRFs. NeRFs [26] enable efficient [27] and photorealistic novel view synthesis by directly optimizing the photometric accuracy of the reconstructed images, in contrast to traditional 3D photogrammetry methods (e.g., for 3D meshes). This provides accurate handling of transparency, reflective materials, and lighting conditions. Ref. [28] employs a NeRF to create a simulator for learning legged robot control policies from RGB images using Reinforcement Learning (RL). While they utilize a specialized camera for data collection, we use the same camera and we incorporate those images into the DA approach. Ref. [29] uses a NeRF for estimation, planning, and control on a drone by querying the NeRF online, but this results in 1000×1000$\times$1000 × higher computation time [^1] Control and estimation in [29] require $6.0$s on a NVIDIA RTX3090 GPU ($10,496$ CUDA cores, $24$ GB VRAM), while ours requires $1.5$ms on a much smaller Jetson TX2 GPU ($256$ CUDA cores, $8$ GB shared RAM). than our policy.
Output feedback RTMPC. MPC [7] solves a constrained optimization problem that uses a model of the system dynamics to plan for actions that satisfy state and actuation constraints. Robust variants of MPC typically assume that the system is subject to additive, bounded process uncertainty (disturbances, model errors), and either a) plan by assuming a worst-case disturbance [30, 31], or b) employs an auxiliary (ancillary) controller, as in RTMPC, that maintains the system within some known distance (cross-section of a tube) of the plan [23]. Output feedback RTMPC [23, 32] in addition accounts for the effects of sensing uncertainty (noise, estimation errors) by increasing the cross-section of the tube. Our method uses an output-feedback RTMPC for data collection but bypasses its computational cost at deployment by learning a NN policy.
@@ III PROBLEM FORMULATION
Our goal is to efficiently train a NN visuomotor policy $\pi_{\theta}$ (student), with parameters $\theta$, that tracks a desired trajectory on a mobile robot (e.g., multirotor). $\pi_{\theta}$ takes as input images, which are needed to extract partial state information (e.g., horizontal position), and other measurements. The trained policy, denoted $\pi_{\hat{\theta}^{*}}$, needs to be robust to uncertainties encountered in the deployment domain $\mathcal{T}$. It is trained using demonstrations from a model-based controller (expert) collected in a source domain $\mathcal{S}$ that presents only a subset of the uncertainties in $\mathcal{T}$.
@@
and it generates deterministic, continuous actions $\mathbf{u}_{t}$ to track a desired $N+1$ steps ($N>0$) trajectory $\mathbf{X}^{\text{des}}_{t}=\{\mathbf{x}^{\text{des}}_{0|t},\dots,\mathbf{x}^{\text{des}}_{N|t}\}$. $\mathbf{o}_{t}=(\mathcal{I}_{t},\mathbf{o}_{\text{other},t})$ are noisy sensor measurements comprised of an image $\mathcal{I}_{t}$ from an onboard camera, and other measurements $\mathbf{o}_{\text{other,t}}$ (e.g., altitude, attitude, velocity).
Expert. Process model: The dynamics of the robot are described by (e.g., via linearization):
@@
where $\mathbf{x}_{t}\in\mathbb{X}\subset\mathfrak{R}^{n_{x}}$ is the state, and $\mathbf{u}_{t}\in\mathbb{U}\subset\mathfrak{R}^{n_{u}}$ the control inputs. The robot is subject to state and input constraints $\mathbb{X}$ and $\mathbb{U}$, assumed convex polytopes containing the origin [23]. $\mathbf{w}_{t}\in\mathbb{W}_{\mathcal{T}}\subset\mathfrak{R}^{n_{x}}$ in (2) captures time-varying additive process uncertainties in $\mathcal{T}$, such as (i) disturbances (e.g., wind for a UAV) (ii) and model changes/errors (e.g., linearization, discretization, and poorly known parameters). $\mathbf{w}_{t}$ is unknown, but the polytopic bounded set $\mathbb{W}_{\mathcal{T}}$ is assumed known[23]. Sensing model: The expert has access to (i) the measurements $\mathbf{o}_{\text{other},t}$, and (ii) a vision-based position estimator $g_{\text{cam}}$ (e.g., via SLAM) that outputs noisy measurements $\mathbf{o}_{\text{pos.xy}}\in\mathfrak{R}^{2}$ of the horizontal position $\bm{p}_{{xy},t}\in\mathfrak{R}^{2}$ of the robot:
@@ IV-A Output feedback robust tube MPC for trajectory tracking
Tube and robust constraints. Process and sensing uncertainties are taken into account by tightening the constraints $\mathbb{X}$, $\mathbb{U}$, obtaining $\bar{\mathbb{X}}$ and $\bar{\mathbb{U}}$ in Eq. 6. The amount by which $\mathbb{X}$, $\mathbb{U}$ are tightened depends on the cross-section of the tube $\mathbb{Z}$, which is computed by observing that the ancillary controller Eq. 7 combined with the system dynamics in Eq. 2 and the observer Eq. 5 constitute a closed-loop system. Sensing uncertainties in $\mathbb{V}_{\mathcal{T}}$ and process noise in $\mathbb{W}_{\mathcal{T}}$ introduce two sources of errors in such system: a) the state estimation error $\bm{\xi}^{\text{est}}_{t}\coloneqq\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}$, and b) the control error $\bm{\xi}^{\text{ctrl}}_{t}\coloneqq\hat{\mathbf{x}}_{t}-\bar{\mathbf{x}}_{t}^{*}$. These errors can be combined in a vector $\bm{\xi}_{t}=[{\bm{\xi}^{\text{est}}_{t}}^{T},{\bm{\xi}^{\text{ctrl}}_{t}}^{T}]^{T}$ whose dynamics evolve according to (see [33, 32]):
@@
$\mathbb{Z}$ (cross-section of a tube) is the set of possible deviations of the true state $\mathbf{x}_{t}$ from the safe reference $\bar{\mathbf{x}}_{t}^{*}$. Figure 2 shows an example of this controller for trajectory tracking on a multirotor, highlighting changes to the reference trajectory to respect state constraint under uncertainties.

### 26 Feb 2024 v2

@@ Abstract
Imitation learning (IL) can train computationally-efficient sensorimotor policies from a resource-intensive Model Predictive Controller (MPC), but it often requires many samples, leading to long training times or limited robustness. To address these issues, we combine IL with a variant of robust MPC that accounts for process and sensing uncertainties, and we design a data augmentation (DA) strategy that enables efficient learning of vision-based policies. The proposed DA method, named Tube-NeRF, leverages Neural Radiance Fields (NeRFs) to generate novel synthetic images, and uses properties of the robust MPC (the tube) to select relevant views and to efficiently compute the corresponding actions. We tailor our approach to the task of localization and trajectory tracking on a multirotor, by learning a visuomotor policy that generates control actions using images from the onboard camera as only source of horizontal position. Numerical evaluations show $80$-fold increase in demonstration efficiency and a $50\%$ reduction in training time over current IL methods. Additionally, our policies successfully transfer to a real multirotor, achieving low tracking errors despite large disturbances, with an onboard inference time of only $1.5$ ms. Video: https://youtu.be/_W5z33ZK1m4
@@ I INTRODUCTION
Imitation learning (IL) [1, 2, 3] has been extensively employed to train sensorimotor neural network (NN) policies for computationally-efficient onboard sensing, planning and control on mobile robots. Training data is commonly provided by a computationally expensive model predictive controller (MPC) [4, 5, 6], acting as expert demonstrators. The resulting policies produce commands from raw sensory inputs, simultaneously bypassing the computational cost of solving large optimization problems in MPC, as well as sensing and localization.
@@
However, one of the fundamental limitations of existing IL methods employed to train sensorimotor policies ( Behavior Cloning (BC)) [2], DAgger [3]) is the overall number of demonstrations that must be collected from MPC. This inefficiency is rooted in training/deployment data distribution mismatches (covariate shift [3]) and is worsened by uncertainties at deployment [7, 8]. Although strategies that match the uncertainties between the training and deployment domains ( Domain Randomization (DR) [9, 10, 4, 5]) improve robustness, they create challenges in data- and computational-efficiency by repeatedly querying a computationally-expensive expert under different disturbances.
@@
Figure 1: Tube-NeRF collects a real-world demonstration using output-feedback tube MPC, a robust MPC which accounts for process and sensing uncertainties through its tube cross-section $\mathbb{Z}$. Then, it generates a Neural Radiance Field (NeRF) of the environment from the collected images $\mathcal{I}_{t}$, and uses the tube’s cross-section to guide the selection of synthetic views $\mathcal{I}_{t}^{+}$ from the NeRF for data augmentation, while corresponding actions are obtained via the ancillary controller, an integral component of the tube MPC framework.
@@
Leveraging high-fidelity simulators on powerful computers in combination with DR avoids demanding data-collection on the real robot, but it introduces sim-to-real gaps that are especially challenging when learning visuomotor policies, i.e., that directly use raw pixels as input. For this reason, sensorimotor policies trained in simulation often leverage as input easy-to-transfer visual abstractions, such as feature tracks [5], depth maps [11, 12], intermediate layers of a convolutional NN (CNN) [13], or a learned latent space [14]. However, all these abstractions discard information that may instead benefit task performance.
@@
Data augmentation (DA) approaches that augment demonstrations with extra images and stabilizing actions improve robustness and sample efficiency of IL [15, 16, 17, 18, 19]. However, they (i) rely on handcrafted heuristics for the selection of extra images and the generation of corresponding actions, (ii) do not explicitly account for the effects of uncertainties when generating the extra data [15, 16, 17, 18, 19], and (iii) often leverage ad-hoc image acquisition setups for DA [16, 17]. As a consequence, their real-world deployment has been mainly focused on tasks in 2D (steering a Dubins car [15, 17]).
@@
This work, as shown in Figure 1, introduces Tube- Neural Radiance Field (NeRF), a novel DA framework for efficient, robust visuomotor policy learning from MPC that overcomes the aforementioned limitations in DA. Building on our prior DA strategy [8] that uses robust tube model predictive controller (RTMPC) for efficiently and systematically generating additional training data for motor control policies (actions from full state), Tube-NeRF enables learning of policies that directly use vision as input, relaxing the constraining assumption in [8] that full-state information is available at deployment.
@@
Our new strategy first collects robust task-demonstrations that account for the effects of process and sensing uncertainties via an output-feedback variant of RTMPC [20]. Then, it employs a photorealistic representation of the environment, based on a NeRF, to generate synthetic novel views for DA, using the tube of the controller to guide the selection of the extra novel views and sensorial inputs, and using the ancillary controller to efficiently compute the corresponding actions. We further enhance DA by organizing collected real-world observations in a database and employing the tube to generate queries. This enables usage of real-world data for DA, reducing the sim-to-real gap when only synthetic images are used. Lastly, we adapt our approach to a multirotor, training a visuomotor policy for robust trajectory tracking and localization using onboard camera images and additional measurements of altitude, orientation, and velocity. The generated policy relies solely on images to obtain information on the robot’s horizontal position, which is a challenging task due to (1) its high speed (up to $3.5$ m/s), (2) varying altitude, (3) aggressive roll/pitch changes, (4) the sparsity of visual features in our flight space, and (5) the presence of a safety net that moves due to the down-wash of the propellers and that produces semi-transparent visual features above the ground.
Contributions: This Letter is an evolution of our previous conference paper [21], where we demonstrated the capabilities of an output-feedback RTMPC-guided DA strategy in simulation. Now, we present a set of algorithmic changes that enabled the first real-world, real-time deployment of the approach. Specifically, different from [21], our new algorithm:
  • Utilizes a NeRF for photo-realistic novel view synthesis for DA, replacing a 3D mesh, and presents a calibration method to integrate the NeRF into the DA framework.
  • Further reduces the sim-to-real gap that, although small, still exists in the synthetic images from the NeRF by modifying the DA strategy to additionally sample available real-world images inside the tube.
  • Accounts for visual changes in the environment by introducing in training randomizations in image space.
  • Includes noisy altitude measurements in the policy’s input.
  • Accounts for camera calibration errors by introducing perturbations in the camera extrinsic during DA.
  • Employs a more capable but computationally efficient visuomotor NN architecture for onboard deployment.
Therefore, this work contributes procedures to:
@@
  • Efficiently (demonstrations, training time) learn of a sensorimotor policy from MPC using a DA strategy grounded in the output feedback RTMPC framework theory, unlike previous DA methods that rely on handcrafted heuristics.
@@
  • Apply our method for tracking and localization on a multirotor via images, altitude, attitude and velocity data.
  • Achieve the first-ever real-time deployment of our approach, demonstrating (in more than $30$ flights) successful agile trajectory tracking with policies learned from a single demonstration that use onboard fisheye images to infer the horizontal position of a multirotor despite aggressive 3D motion, and subject to a variety of sensing and dynamics disturbances. Our policy has an average inference time of only $1.5$ ms onboard a small GPU (Nvidia Jetson TX2) and is deployed at $200$ Hz. The policy is trained using a single demonstration.
@@ II RELATED WORKS
Method Domain of training data Policy directly uses images No special data collection equipment Demo.-efficient Explicitly accounts for uncertainties Avoids Hand-Crafted Heuristics for Data Augm. Real-world deployment (2D/3D, Domain)
@@
Novel view synthesis with NeRFs. NeRFs [23] enable efficient [24] and photorealistic novel view synthesis by directly optimizing the photometric accuracy of the reconstructed images, in contrast to traditional 3D photogrammetry methods (e.g., for 3D meshes). This provides accurate handling of transparency, reflective materials, and lighting conditions. Ref. [25] employs a NeRF to create a simulator for learning legged robot control policies from RGB images using Reinforcement Learning (RL), using a specialized camera for data collection, unlike our work. Ref. [26] uses a NeRF for estimation, planning, and control on a drone by querying the NeRF online, but this results in 1000×1000$\times$1000 × higher computation time [^1] Control and estimation in [26] require $6.0$s on a NVIDIA RTX3090 GPU ($10,496$ CUDA cores, $24$ GB VRAM), while ours requires $1.5$ms on a much smaller Jetson TX2 GPU ($256$ CUDA cores, $8$ GB shared RAM). than our policy.
Output feedback RTMPC. MPC [27] solves a constrained optimization problem that uses a model of the system dynamics to plan for actions that satisfy state and actuation constraints. RTMPC assumes that the system is subject to additive, bounded process uncertainty (disturbances, model errors) and employs an auxiliary (ancillary) controller that maintains the system within some known distance (cross-section of a tube) of the plan [20]. Output feedback RTMPC [20, 28] in addition accounts for the effects of sensing uncertainty (noise, estimation errors) by increasing the cross-section of the tube. Our method uses an output-feedback RTMPC for data collection but bypasses its computational cost at deployment by learning a NN policy.
@@ III PROBLEM FORMULATION
Our goal is to efficiently train a NN visuomotor policy $\pi_{\theta}$ (student), with parameters $\theta$, that tracks a desired trajectory on a mobile robot (multirotor). $\pi_{\theta}$ takes as input images, which are needed to extract partial state information (horizontal position, in our evaluation), and other measurements. The trained policy, denoted $\pi_{\hat{\theta}^{*}}$, needs to be robust to uncertainties encountered in the deployment domain $\mathcal{T}$. It is trained using demonstrations from a model-based controller (expert) collected in a source domain $\mathcal{S}$ that presents only a subset of the uncertainties in $\mathcal{T}$.
@@
and it generates deterministic, continuous actions $\mathbf{u}_{t}$ to track a desired $N+1$ steps ($N>0$) trajectory $\mathbf{X}^{\text{des}}_{t}=\{\mathbf{x}^{\text{des}}_{0|t},\dots,\mathbf{x}^{\text{des}}_{N|t}\}$. $\mathbf{o}_{t}=(\mathcal{I}_{t},\mathbf{o}_{\text{other},t})$ are noisy sensor measurements comprised of an image $\mathcal{I}_{t}$ from an onboard camera, and other measurements $\mathbf{o}_{\text{other,t}}$ (altitude, attitude, velocity, in our evaluation).
Expert. Process model: The considered robot dynamics are:
@@
where $\mathbf{x}_{t}\in\mathbb{X}\subset\mathfrak{R}^{n_{x}}$ is the state, and $\mathbf{u}_{t}\in\mathbb{U}\subset\mathfrak{R}^{n_{u}}$ the control inputs. The robot is subject to state and input constraints $\mathbb{X}$ and $\mathbb{U}$, assumed convex polytopes containing the origin [20]. $\mathbf{w}_{t}\in\mathbb{W}_{\mathcal{T}}\subset\mathfrak{R}^{n_{x}}$ in (2) captures time-varying additive process uncertainties in $\mathcal{T}$, such as (i) disturbances (wind/payloads for a UAV) (ii) and model changes/errors (linearization errors and poorly known parameters). $\mathbf{w}_{t}$ is unknown, but the polytopic bounded set $\mathbb{W}_{\mathcal{T}}$ is assumed known[20]. Sensing model: The expert has access to (i) the measurements $\mathbf{o}_{\text{other},t}$, and (ii) a vision-based position estimator $g_{\text{cam}}$ that outputs noisy measurements $\mathbf{o}_{\text{pos.xy}}\in\mathfrak{R}^{2}$ of the horizontal position $\bm{p}_{{xy},t}\in\mathfrak{R}^{2}$ of the robot:
@@ IV-A Output feedback robust tube MPC for trajectory tracking
Tube and robust constraints. Process and sensing uncertainties are taken into account by tightening the constraints $\mathbb{X}$, $\mathbb{U}$, obtaining $\bar{\mathbb{X}}$ and $\bar{\mathbb{U}}$ in Eq. 6. The amount by which $\mathbb{X}$, $\mathbb{U}$ are tightened depends on the cross-section of the tube $\mathbb{Z}$, which is computed (see [20]) by considering the closed-loop system formed by the ancillary controller Eq. 7, the nominal dynamics in Eq. 2 and the observer Eq. 5. Sensing uncertainties in $\mathbb{V}_{\mathcal{T}}$ and process noise in $\mathbb{W}_{\mathcal{T}}$ introduce two sources of errors in such system: a) the state estimation error $\bm{\xi}^{\text{est}}_{t}\coloneqq\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}$, and b) the control error $\bm{\xi}^{\text{ctrl}}_{t}\coloneqq\hat{\mathbf{x}}_{t}-\bar{\mathbf{x}}_{t}^{*}$. These errors can be combined in a vector $\bm{\xi}_{t}=[{\bm{\xi}^{\text{est}}_{t}}^{T},{\bm{\xi}^{\text{ctrl}}_{t}}^{T}]^{T}$ whose dynamics evolve according to (see [28]):
@@
$\mathbb{Z}$ (cross-section of a tube) is the set of possible deviations of the true state $\mathbf{x}_{t}$ from the safe reference $\bar{\mathbf{x}}_{t}^{*}$. Computing $\mathbb{S}$. While accurately computing the minimal RPI set $\mathbb{S}$ for high-dimensional systems can be challenging [28], for simplicity, we efficiently obtain $\mathbb{S}$ from $\mathbb{W}_{\mathcal{T}}$ and $\mathbb{V}_{\mathcal{T}}$ via Monte Carlo simulations of Eq. 8, uniformly sampling instances of the uncertainties, and by computing an outer axis-aligned bounding box of the trajectories of $\bm{\xi}$. In addition, we treat linearization errors and future changes in the reference trajectory as an additional source of uncertainty, computing the tube based on an increased process uncertainty prior $\bar{\mathbb{W}}_{\mathcal{T}}$, and numerically validating that the resulting expert is robust. While this procedure is approximate, it was found computationally tractable and useful at estimating tubes with an adequate level of conservativeness. Figure 2 shows an example of this controller for trajectory tracking on a multirotor, highlighting changes to the reference trajectory to respect state constraint under uncertainties.